OpenStudy (anonymous):

what is the root of[9x-8=x^{2}

OpenStudy (anonymous):

positive 8 and 1

OpenStudy (anonymous):

You subtract the 9x and add the 8 to the other side withx^2

OpenStudy (anonymous):

then, you factor that to get (x-1) and (x-8)

OpenStudy (anonymous):

after that, set them to equal zero like so: x-1= 0 and x-8= 0

OpenStudy (anonymous):

what is the roots for these 3problems x^{2}-7x-4=10 x^{2}-20x+99=0 x^{2}=x-72=0

OpenStudy (anonymous):

for the first one add 10 to the opposite side.

OpenStudy (anonymous):

you notice that these do not have any possible multiples to equal to 14

OpenStudy (anonymous):

so what you do is, take the equation : b+ or - the square root of b^2 -4(a)(c)

OpenStudy (anonymous):

that will give you -7+or- the square root of -7^2-4(1)(14)

OpenStudy (anonymous):

that will be -7+or- the square root of 7

OpenStudy (anonymous):

the final answer for that one is -7+or- the square root of 7 over 2. I forgot to tell you that the equation I posted ealier:b+ or - the square root of b^2 -4(a)(c) - is divided by two. You use this equation when the mulitples of the third term in a problem cannot give you the middle term when added or subtracted